# Experience Replay with Random Reshuffling

### v1

@@ Appendix A Sampling methods of experience replay in popular RL code bases
TorchRL (Bou et al., 2023) WR [^19] https://github.com/pytorch/rl/blob/21c4d87c7ba5592a731757411468d080b7173b5f/torchrl/data/replay_buffers/storages.py#L159202020Additionally, TorchRL also supports SamplerWithoutReplacement sampler, which resembles our RR-C. However, it resets its list of shuffled indices every time the storage is expanded, thus making it unsuitable for a replay buffer of a growing size, which is typical in off-policy online RL training. https://github.com/pytorch/rl/blob/21c4d87c7ba5592a731757411468d080b7173b5f/torchrl/data/replay_buffers/samplers.py#L162 WR [^21] https://github.com/pytorch/rl/blob/21c4d87c7ba5592a731757411468d080b7173b5f/torchrl/data/replay_buffers/samplers.py#L497
Appendix B Example trajectories of RR-M
@@
We provide example trajectories of RR-M to illustrate its behavior. We use the simple example from Section 3.3, where three transitions with priorities $[1,0.5,2]$ are in the buffer, and we sequentially sample 7 transitions. Two trajectories simulated using different random seeds are shown in Table 4. Both trajectories sample transitions in different orders, but eventually they sample the first transition twice, the second once, and the third four times, which is the expected behavior of RR-M.
@@
Appendix C On the computational efficiency of RR-M
@@
C.1 Computational cost of RR-M
RR-M tracks actual and expected sample counts, updating them every time a minibatch is sampled Actual counts are incremented only for sampled transitions, requiring $O(b)$ time, where $b$ is the minibatch size. In contrast, expected counts are updated for all transitions in the buffer, so it takes $O(n)$ time. Identifying oversampled transitions involves comparing the actual and expected counts, which also requires $O(n)$ time.
@@ C.1 Computational cost of RR-M
C.2 Efficient implementation of RR-M
Although updating expected counts and the identification of oversampled transitions each require $O(n)$ time, both tasks can be efficiently parallelized on GPUs. We implement two sum trees, actual and expected sample counts, and a binary mask as GPU arrays using CuPy (Okuta et al., 2017), so we can update them efficiently. Our implementation is included in the supplementary materials.
@@ C.2 Efficient implementation of RR-M
Appendix D Experimental details
D.1 Additional simulation results
We provide additional results of experience replay simulations with different parameters in Figure 4, Figure 5, and Figure 6.
Figure 4: Distributions of sample counts in experience replay simulations with a different set of parameters from Figure 3: the minibatch size is 8. Format follows Figure 3.
Figure 5: Distributions of sample counts in experience replay simulations with a different set of parameters from Figure 3: the total timesteps is 1000, the capacity of the buffer size is 200, the size at which replay starts is 100, and $p_{t}=(t\bmod 250)+50$. Format follows Figure 3.
Figure 6: Distributions of sample counts in experience replay simulations with a different set of parameters from Figure 3: the total timesteps is 1000, the capacity of the buffer size is 200, the size at which replay starts is 100, the minibatch size is 8, and $p_{t}=(t\bmod 250)+50$. Format follows Figure 3.
D.2 Experimental setup
@@
D.3 Final performance in details
We provide the final performance of C51, DQN, and DDQN+LAP with different sampling methods in the same way as Table 1 except that we report the mean $\pm$ standard deviations (ddof=1) in Table 5, Table 6, Table 7, Table 8, and Table 9.
Table 5: Final performance of C51 with with-replacement sampling (WR) and RR-C. Format follows Table 1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@ D.3 Final performance in details
Table 6: Final performance of DQN with with-replacement sampling (WR) and RR-C. Format follows Table 1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
Table 7: Final performance of C51 with with-replacement sampling (WR) and within-minibatch without-replacement sampling (WOR). Format follows Table 1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
Table 8: Final performance of DDQN+LAP with with-replacement sampling (WR) and RR-M. Format follows Table 1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
Table 9: Final performance of DDQN+LAP with stratified sampling (ST) and RR-M+ST. Format follows Table 1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
D.4 Learning curves
We provide the learning curves of C51, DQN, and DDQN+LAP with different sampling methods in Figure 7, Figure 8, Figure 9, Figure 10, and Figure 11.
@@ D.4 Learning curves
Figure 11: Learning curves of DDQN+LAP with stratified sampling (ST) and RR-M+ST. The 500-episode rolling mean of episodic returns is plotted for each of 10 random seeds.

### v2

@@ Appendix A Sampling methods of experience replay in popular RL code bases
TorchRL (Bou et al., 2023) WR [^19] https://github.com/pytorch/rl/blob/21c4d87c7ba5592a731757411468d080b7173b5f/torchrl/data/replay_buffers/storages.py#L159202020Additionally, TorchRL also supports SamplerWithoutReplacement sampler, which resembles our $R$-C. However, it resets its list of shuffled indices every time the storage is expanded, thus making it unsuitable for a replay buffer of a growing size, which is typical in off-policy online RL training. https://github.com/pytorch/rl/blob/21c4d87c7ba5592a731757411468d080b7173b5f/torchrl/data/replay_buffers/samplers.py#L162 WR [^21] https://github.com/pytorch/rl/blob/21c4d87c7ba5592a731757411468d080b7173b5f/torchrl/data/replay_buffers/samplers.py#L497
Appendix B Theoretical results
In this section, we give theoretical results with proofs on our proposed sampling methods, RR-C and RR-M.
B.1 Notations
We denote by $C$ the capacity of the replay buffer and by $B$ the minibatch size. For every timestep $t=0,1,\dots$ we add a new transition $T_{t}$ to the buffer, and, if the buffer contains at least $R$ transitions, sample a minibatch of size $B$ from the buffer. $B\leq R\leq C$ always holds. For prioritized settings, let $p_{i}$ be the priority of transition $T_{i}$ in the buffer.
The number of times transition $T_{i}$ is sampled up to timestep $t=k$ is a random variable, denoted $X_{i,k}$, where randomness comes from the sampling procedure. Since $T_{i}$ is added to the replay buffer at timestep $t=i$ and dropped at timestep $t=i+C$, $X_{i,k}=0$ for all $k<i$ and $X_{i,k}=X_{i,i+C-1}$ for all $k\geq i+C-1$. Let $t_{\mathrm{start}}(i)\coloneq\max\{i,\;R-1\}$ and $t_{\mathrm{end}}(i,k)\coloneq\min\{k,\;i+C-1\}$. The number of timesteps on which minibatches are actually drawn while $T_{i}$ is present is $N_{i,k}\coloneq\max\{0,\;t_{\mathrm{end}}(i,k)-t_{\mathrm{start}}(i)+1\}$, and we define $S_{i,k}\coloneq B\cdot N_{i,k}$. In particular, when $i\geq C$ we have $t_{\mathrm{start}}(i)=i$ and $t_{\mathrm{end}}(i,k)=\min\{k,i+C-1\}$, so $S_{i,k}=B\cdot\min(k-i+1,C)$, which is the expression we will use in the proofs below.
We represent which sampling method is used by a superscript of a random variable, so $X_{i,k}^{\text{UER}}$, $X_{i,k}^{\text{PER}}$, $X_{i,k}^{\text{RR-C}}$, and $X_{i,k}^{\text{RR-M}}$ denote the number of times transition $T_{i}$ is sampled up to timestep $t=k$ when using uniform experience replay with with-replacement sampling (UER), prioritized experience replay with with-replacement sampling (PER), RR-C, and RR-M, respectively.
B.2 Bias of RR-C
First, we show that RR-C can be biased when the skipping mechanism in the index buffer (rr_buffer in figure˜1) is activated. {theorem} There exists a configuration such that $\mathbb{E}\left[X_{i,k}^{\text{RR-C}}\right]\neq\mathbb{E}\left[X_{i,k}^{\text{UER}}\right]$ for some $(i,k)$.
Proof.
Let $B=R=1,C=2$ and $(i,k)=(0,1)$.
In uniform experience replay, $T_{0}$ is sampled at $k=0$ with probability $1$ and at $k=1$ with probability $1/2$, so $\mathbb{E}\left[X_{0,1}^{\text{UER}}\right]=1+1/2=3/2$.
In RR-C, at $t=0$, the index buffer is initialized as either $[0,1]$ or $[1,0]$ with probability $1/2$ each.
  • (a) If the index buffer is $[0,1]$, then $T_{0}$ is sampled at $t=0$, leaving the index buffer as $[1]$. At $t=1$, $T_{1}$ is sampled.
  • (b) If the index buffer is $[1,0]$, then $1$ is skipped and $T_{0}$ is sampled at $t=0$. Now that the index buffer is empty, the index buffer is regenerated as either $[0,1]$ or $[1,0]$ with probability $1/2$ each.
    • (b1) If the regenerated index buffer is $[0,1]$, then $T_{0}$ is sampled at $t=1$.
    • (b2) If the regenerated index buffer is $[1,0]$, then $T_{1}$ is sampled at $t=1$.
(a), (b1), and (b2) occur with probability $1/2$, $1/4$, and $1/4$, respectively. Thus, $\mathbb{E}\left[X_{0,1}^{\text{RR-C}}\right]=1\cdot(1/2+1/4)+2\cdot(1/4)=5/4\neq 3/2$. ∎
However, RR-C is unbiased for sufficiently large $i$ so that the skipping mechanism is never activated. {theorem}[store*=rrc_bias] $\mathbb{E}\left[X_{i,k}^{\text{RR-C}}\right]=\mathbb{E}\left[X_{i,k}^{\text{UER}}\right]$ holds for all $(i,k)$ for sufficiently large $i$.
Proof.
First, we note that for timestep $t\geq C$, the replay buffer is always full. To exhaust the whole index buffer of size $C$, we need $\lceil C/B\rceil$ minibatches, i.e. $\lceil C/B\rceil$ timesteps. Thus, there must be a timestep $t^{*}$ in $C\leq t^{*}\leq C+\lceil C/B\rceil$ such that the index buffer is regenerated at $t^{*}$ and there is no skipping thereafter. For any $i>t^{\star}$ and any $k\geq i$, all minibatch draws that can include $T_{i}$ occur at timesteps $t\in[i,\min\{k,i+C-1\}]$ with $t>t^{\star}$, so the skipping mechanism never affects $X^{\text{RR-C}}_{i,k}$.
In uniform experience replay, as long as the buffer is full, i.e., $i\geq C$, when we sample a transition, the probability of it being transition $T_{i}$ is $1/C$ as long as it is in the buffer. Thus, for $i\geq C$, $\mathbb{E}\left[X_{i,k}^{\text{UER}}\right]=S_{i,k}\cdot(1/C)$ for all $k\geq i$.
To analyze the case of RR-C, we decompose $S_{i,k}$ as $S_{i,k}=l+mC+n$, where $m$ is a number of times the index buffer is fully exhausted, $l$ is the number of samples taken before the first exhaustion, and $n$ is the number of samples taken after the last exhaustion. We can consider the index buffer is uniformly shuffled when $i$ is sufficiently large, as argued above. Thus, the expected number of times $T_{i}$ appears in the first $l$ draws is $l/C$, and in the last $n$ draws is $n/C$. Each of the $m$ full exhaustions samples $T_{i}$ exactly once. Therefore, $\mathbb{E}\left[X_{i,k}^{\text{RR-C}}\right]=l/C+m+n/C=(l+m*C+n)/C=S_{i,k}/C=\mathbb{E}\left[X_{i,k}^{\text{UER}}\right]$. ∎
Thus, RR-C can be biased only early in training. We have not quantified the bias further, but our numerical simulations in figures˜3, 4, 5 and 6 suggest that the bias is negligible in practice.
B.3 Variance of RR-C
Next, we show that RR-C achieves lower variance in the number of times each transition is sampled compared to uniform experience replay under practical conditions.
{theorem}
[store*=rrc_var] $\mathrm{Var}\left[X_{i,k}^{\text{RR-C}}\right]\leq\mathrm{Var}\left[X_{i,k}^{\text{UER}}\right]$ holds for all $(i,k)$ such that $i\leq k$ with sufficiently large $i$. The equality holds only under a few specific conditions.
Proof.
As in the proof of unbiasedness of RR-C, we consider the case where $i$ is sufficiently large so that there is no skipping in the index buffer.
In uniform experience replay, whether each sample is $T_{i}$ is an independent Bernoulli trial with success probability $1/C$. Since we draw $S=B\cdot\min(k-i+1,C)$ samples from $t=i$ to $t=k$, we obtain $\mathrm{Var}\left[X_{i,k}^{\text{UER}}\right]=S_{i,k}\cdot(1/C)\cdot(1-1/C)$.
In RR-C, we again use the decomposition $S_{i,k}=l+mC+n$ from the proof of unbiasedness of RR-C. The variance of $X_{i,k}^{\text{RR-C}}$ comes from the first $l$ samples and the last $n$ samples, as the $m$ full exhaustions always include $T_{i}$. The event that $T_{i}$ appears in the first $l$ draws and the event that it appears in the last $n$ draws correspond to different epochs (different independent permutations of the index buffer), so these two Bernoulli variables are independent. Thus, $\mathrm{Var}\left[X_{i,k}^{\text{RR-C}}\right]=(l/C)\cdot(1-l/C)+(n/C)\cdot(1-n/C)$.
Now we compare the two variances.
$\mathrm{Var}\left[X_{i,k}^{\text{UER}}\right]-\mathrm{Var}\left[X_{i,k}^{\text{RR-C}}\right]$ $=S_{i,k}\cdot(1/C)\cdot(1-1/C)-(l/C)\cdot(1-l/C)-(n/C)\cdot(1-n/C)$
$=\frac{l(l-1)+n(n-1)+mC(C-1)}{C^{2}}$
Since $0\leq l<C,0\leq m,0\leq n<C$, we obtain $\mathrm{Var}\left[X_{i,k}^{\text{RR-C}}\right]\leq\mathrm{Var}\left[X_{i,k}^{\text{UER}}\right]$, where the equality holds only under either of the following conditions: (a) $C=1$ or (b) $l\in\{0,1\},m=0,n\in\{0,1\}$. ∎
The condition (a) is trivial as the buffer can contain only one transition. The condition (b) corresponds to the case where (b1) $i=k$ or (b2) $i+1=k$ and the index buffer is regenerated immediately after $t=i$. As we are interested in the variance over multiple timesteps, the condition (b) is not relevant in practice.
Our numerical simulations in figures˜3, 4, 5 and 6 support this theoretical result, showing that RR-C achieves lower variance throughout training. While our analysis does not cover early timesteps, the variance for the transitions from early timesteps is also reduced compared to uniform experience replay in our simulations.
B.4 Bias of RR-M
Similarly to RR-C, we show that RR-M can be biased. {theorem} There exists a configuration such that $\mathbb{E}\left[X_{i,k}^{\text{RR-M}}\right]\neq\mathbb{E}\left[X_{i,k}^{\text{PER}}\right]$ for some $(i,k)$.
Proof.
Let $B=1,R=2,C=3,p_{0}=0.6,p_{1}=0.4,p_{2}=0.0$ and $(i,k)=(0,2)$.
In prioritized experience replay, $T_{0}$ is sampled at $t=1,2$ with probability $0.6$ so $\mathbb{E}\left[X_{0,2}^{\text{PER}}\right]=0.6+0.6=1.2$.
In RR-M, at $t=1$, $T_{0}$ is sampled with probability $0.6$ and $T_{1}$ with $0.4$.
  • (a) If $T_{0}$ is sampled at $t=1$, then the expected and the actual counts are $[0.6,0.4],[1,0]$, respectively. At $t=2$, since the actual count of $T_{0}$ is greater than the expected count, $T_{0}$ is masked out and $T_{1}$ is sampled.
  • (b) If $T_{1}$ is sampled at $t=1$, then the expected and the actual counts are $[0.6,0.4],[0,1]$, respectively. At $t=2$, since the actual count of $T_{1}$ is greater than the expected count, $T_{1}$ is masked out and $T_{0}$ is sampled.
Thus, $\mathbb{E}\left[X_{0,2}^{\text{RR-M}}\right]=0.6\cdot 1+0.4\cdot 1=1\neq 1.2$. ∎
Unlike RR-C, this bias can be nonzero even for large $i$, e.g., when $p_{i}=0.6,p_{i+1}=0.4$ and other priorities are zero, the same argument as above applies for all $(i,i+1)$.
Due to the dynamic masking mechanism of RR-M, it is difficult to analyze its bias accurately. Below, we consider a simplified setting where we stop adding a new transition somewhere after $t=i$ so that $T_{i}$ is forever kept in the buffer. We also assume $B=1$ for simplicity.
Lemma 1.
$1-C<X_{i,k}^{\text{RR-M}}-\mathbb{E}\left[X_{i,k}^{\text{PER}}\right]<1$ holds for all $(i,k)$ under a simplified setting.
Proof.
Define $A_{i}(t)$ and $E_{i}(t)$ as the actual and expected counts of $T_{i}$ at timestep $t$ after increment, respectively. Then, we have $X_{i,k}^{\text{RR-M}}=A_{i}(k)$ and $\mathbb{E}\left[X_{i,k}^{\text{PER}}\right]=E_{i}(k)$.
By construction of RR-M, $T_{i}$ is never sampled at $t$ when $A_{i}(t-1)>E_{i}(t-1)$, which prevents $A_{i}(t)$ to be incremented further. Thus, we have $A_{i}(t)<E_{i}(t)+1$ for all $t$, which implies the upper bound, $X_{i,k}^{\text{RR-M}}-\mathbb{E}\left[X_{i,k}^{\text{PER}}\right]<1$.
Also, $\sum_{i}A_{i}(t)=\sum_{i}E_{i}(t)$ must hold for any $t$, since both actual and expected counts are incremented by exactly one in total at each timestep. Using this fact and $A_{i}(t)<E_{i}(t)+1$, we have the lower bound $1-C<X_{i,k}^{\text{RR-M}}-\mathbb{E}\left[X_{i,k}^{\text{PER}}\right]$
Thus, under the simplified setting, the empirical sampling frequency of RR-M converges to that of PER, in the sense of vanishing relative error. {theorem} $\lim_{k\to\infty}\frac{X_{i,k}^{\text{RR-M}}}{\mathbb{E}\left[X_{i,k}^{\text{PER}}\right]}=1$ holds almost surely for all $i$ under a simplified setting.
Proof.
From lemma˜1, we have $\lvert A_{i}(t)-E_{i}(t)\rvert<K$, where $K$ is a constant independent of $t$. Dividing both sides by $E_{i}(t)$ and taking the limit $t\to\infty$, we obtain the result since $E_{i}(t)\to\infty$ as $t\to\infty$. ∎
While we only have this asymptotic result in a simplified setting, our numerical simulations in figures˜3, 4, 5 and 6 suggest that the bias is negligible in practice.
B.5 Variance of RR-M
Using the simplified setting as in the previous subsection, we can also bound the variance of RR-M.
Lemma 2.
$\mathrm{Var}\left[X_{i,k}^{\text{RR-M}}\right]\leq\frac{1}{4}C^{2}$ holds for all $(i,k)$ under a simplified setting.
Proof.
From lemma˜1, we have $X_{i,k}^{\text{RR-M}}\in\left[1-C-\mathbb{E}\left[X_{i,k}^{\text{PER}}\right],1-\mathbb{E}\left[X_{i,k}^{\text{PER}}\right]\right]$. We apply Popoviciu’s inequality on variances to $X_{i,k}^{\text{RR-M}}$. ∎
Note that this variance bound is independent of $k$, meaning that the variance does not grow with the number of samples.
We can now compare the variance of RR-M with that of prioritized experience replay. {theorem}[store*=rrm_var] $\mathrm{Var}\left[X_{i,k}^{\text{RR-M}}\right]<\mathrm{Var}\left[X_{i,k}^{\text{PER}}\right]$ holds for all $(i,k)$ with sufficiently large $k$ under a simplified setting.
Proof.
It follows from lemma˜2 and the fact that the variance of prioritized experience replay grows linearly with the number of samples as $\mathrm{Var}\left[X_{i,k}^{\text{PER}}\right]=(k-i+1)\cdot p_{i}\cdot(1-p_{i})$. ∎
Our numerical simulations in figures˜3, 4, 5 and 6 suggest that the variance of RR-M is lower than that of prioritized experience replay throughout training for more realistic settings.
Appendix C Example trajectories of RR-M
@@
We provide example trajectories of RR-M to illustrate its behavior. We use the simple example from section˜3.3, where three transitions with priorities $[1,0.5,2]$ are in the buffer, and we sequentially sample 7 transitions. Two trajectories simulated using different random seeds are shown in table˜4. Both trajectories sample transitions in different orders, but eventually they sample the first transition twice, the second once, and the third four times, which is the expected behavior of RR-M.
@@
Appendix D On the computational efficiency of RR-M
@@
D.1 Computational cost of RR-M
RR-M tracks actual and expected sample counts, updating them every time a minibatch is sampled. Actual counts are incremented only for sampled transitions, requiring $O(b)$ time, where $b$ is the minibatch size. In contrast, expected counts are updated for all transitions in the buffer, so it takes $O(n)$ time. Identifying oversampled transitions involves comparing the actual and expected counts, which also requires $O(n)$ time.
@@ D.1 Computational cost of RR-M
D.2 Efficient implementation of RR-M
Although updating expected counts and the identification of oversampled transitions each require $O(n)$ time, both tasks can be efficiently parallelized on GPUs. We implement two sum trees, actual and expected sample counts, and a binary mask as GPU arrays using CuPy (Okuta et al., 2017), so we can update them efficiently. Our implementation is available at https://github.com/pfnet-research/errr.
@@ D.2 Efficient implementation of RR-M
Appendix E Experimental details
E.1 Additional simulation results
We provide additional results of experience replay simulations with different parameters in figures˜4, 5 and 6.
Figure 4: Distributions of sample counts in experience replay simulations with a different set of parameters from figure˜3: the minibatch size is 8. Format follows figure˜3.
Figure 5: Distributions of sample counts in experience replay simulations with a different set of parameters from figure˜3: the total timesteps is 1000, the capacity of the buffer size is 200, the size at which replay starts is 100, and $p_{t}=(t\bmod 250)+50$. Format follows figure˜3.
Figure 6: Distributions of sample counts in experience replay simulations with a different set of parameters from figure˜3: the total timesteps is 1000, the capacity of the buffer size is 200, the size at which replay starts is 100, the minibatch size is 8, and $p_{t}=(t\bmod 250)+50$. Format follows figure˜3.
E.2 Experimental setup
@@
E.3 Final performance in details
We provide the final performance of C51, DQN, and DDQN+LAP with different sampling methods in the same way as table˜1 except that we report the mean $\pm$ standard deviations (ddof=1) in tables˜5, 6, 7, 8 and 9.
Table 5: Final performance of C51 with with-replacement sampling (WR) and $R$-C. Format follows table˜1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@ E.3 Final performance in details
Table 6: Final performance of DQN with with-replacement sampling (WR) and $R$-C. Format follows table˜1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
Table 7: Final performance of C51 with with-replacement sampling (WR) and within-minibatch without-replacement sampling (WOR). Format follows table˜1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
Table 8: Final performance of DDQN+LAP with with-replacement sampling (WR) and RR-M. Format follows table˜1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
Table 9: Final performance of DDQN+LAP with stratified sampling (ST) and $R$-M+ST. Format follows table˜1. Mean $\pm$ standard deviation (ddof=1) is reported.
@@
E.4 Learning curves
We provide the learning curves of C51, DQN, and DDQN+LAP with different sampling methods in figures˜7, 8, 9, 10 and 11.
@@ E.4 Learning curves
Figure 11: Learning curves of DDQN+LAP with stratified sampling (ST) and $R$-M+ST. The 500-episode rolling mean of episodic returns is plotted for each of 10 random seeds.